(6,1) (5,4) graph a parallel and perpendicular line and equation in any form.
Answer by richwmiller(17219)

m$=3/-1
m=3/-1=-3
m=3/-1
b=-1*3/-1*x1+1*y1
b=-1*3/-1*6+1*1=19
y= 3/-1x+19
y= -3x+19 is a line though those points
slope is -3
so y= -3x is a parallel line
perpendicular line will have a slope of 1/3
y=x/3 is perpendicular